A certain mass of a gas at standard temperature and pressure (STP) is expanded to 3 times its
o rigin a l v o lu m e un d e r ad iab a tic co n d itio n s. C a lc u la te th e re su ltin g te m p e ra tu re a n d pressure. (Take y = 1.4).
1
Expert's answer
2020-06-09T13:22:02-0400

For adiabatic processes "pV^{\gamma}=\text{const}" . Here "\gamma=1.4" , therefore for initial and final conditions

"p_1V_1^{\gamma} = p_2V_2^{\gamma}" . We know that "V_2 = 3V_1," therefore

"p_2 = p_1 \left(\dfrac{V_1}{V_2} \right)^{\gamma} = p_1 \left(\dfrac13 \right)^{\gamma} ." Standard temperature and pressure imply the temperature "T=273.15\,\mathrm{K}, p = 10^5\,\mathrm{Pa}." So "p_1 = 10^5\,\mathrm{Pa}" and

"p_2 = 10^5\,\mathrm{Pa}\cdot \left(\dfrac13 \right)^{1.4} \approx 2.15 \cdot10^4\,\mathrm{Pa}."

According to the ideal gas law,

"p_1V_1 = \nu RT_1, \; p_2V_2 = \nu RT_2," so

"\dfrac{p_2V_2}{p_1V_1} = \dfrac{T_2}{T_1}, \;\; T_2 = T_1 \cdot \dfrac{p_2V_2}{p_1V_1} = 273.15\cdot \dfrac{2.15\cdot10^4\,\mathrm{Pa}\cdot3}{10^5\,\mathrm{Pa}} = 176\,\mathrm{K}."